Raven Applies overhead based on direct labor hours. The variable overhead standard is 2 hours at $11 per hour . During July, Raven spent $116,700 for variable overhead 8,890 labor hours were used to produce 4,700 units. What is variable overhead efficiency variance?

Venus Company applies overhead based on direct labor hours. The variable overhead standard is 10 hours at $3.50 per hour. During October, Venus company spent $157,600 for variable overhead. 47,440 labor hours were used to produce 4,880 units. What is the over or under applied variable overhead ?
Answer #1

1. Variable overhead efficiency variance

= (SH - AH) *SR

= (4700*2 - 8890)*11

= (9400 - 8890)*11

= 5610 Favourable